Show your work AND explain your reasoning using complete English sentences. Explanations must directly follow the solution for each part of the problem.

As of 2017, the following are the ages, in chronological order, at which the last 30 U.S. Presidents were inaugurated:

70 47 54 46 64 69 52 61 56 55 43 62 60 51 54 51 55 56 51 42 54 55 55 55 47 50 49 54 46 56

a. Organize the data by creating a frequency distribution using the following classes: 40-44, 45-49, 50-54, 55-59, 60-64, 65-69, and 70- 74

b. How many of the last 30 U.S. Presidents were inaugurated while in their 50’s?

c. What percentage of the last 30 U.S. Presidents were inaugurated at the age of 60 or later? (Round to 1 decimal place.)
Answer #1

a)

class frequency
40-44 2
45-49 5
50-54 9
55-59 8
60-64 4
65-69 1
70-74 1

b)

number of last 30 U.S. Presidents were inaugurated while in their 50’s =9+8=17

c)

percentage of the last 30 U.S. Presidents were inaugurated at the age of 60 or later =(4+1+1)*100/30

=20 %